OpenStudy (anonymous):

0<=t<=5. find when t(cost)>0

sam (.sam.):

You have (t)(cost)>0 t>0 and cos(t)>0 t>arccos(0) t>90 t>270

OpenStudy (anonymous):

howd you go from 90 to 270? @.Sam.

OpenStudy (anonymous):

you should look for the domain of cost for which cos t >0

OpenStudy (anonymous):

ie from \[-\pi/2 -\pi /2\]

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

ya

sam (.sam.):

because cos(0)=90, then the 270 comes from  so, 360-90= 270